Using polynomials times monomials how do i set up the steps in this problem 4(2n + 3) - 3(n - 1) = 0
Answer by Edwin McCravy(6940)

Using polynomials times monomials how do i set up the
steps in this problem

 4(2n + 3) - 3(n - 1) = 0

Get rid of the parentheses using the distributive 
property:

Multiply the 4 times the 2n which gives 8n
Multiply the 4 times the +3 which gives +12
Multiply the -3 times the n which gives -3n
Multiply the -3 times the -1 which gives +3

So the equation is now free of parentheses:

8n + 12 - 3n + 3 = 0

Combine the 8n and the -3n to get 5n
Combine the +12 and the +3 to get +15

5n + 15 = 0

Add -15 to both sides

5n + 15 = 0
    -15   -15
-------------
5n      = -15

Divide both sides by coefficient 5

5n    -15
-- = -----
 5     5

Cancel the 5's on the left, getting
n, and divide -15 by 5 on the right,
getting -3

 n = -3

Edwin